OpenStudy (anonymous):

Simplify With All Positive Exponents : (u^-2/ w^3)(-2u^2v^3w)^-3

OpenStudy (anonymous):

hi mirene do you have the selection answers? (u^-2/ w^3)(-2u^2v^3 w)^-3 =[1/(u^2 w^3)][1/(-2u^2v^3 w)^3] =[1/(u^2 w^3)][-1/(8u^6v^9 w^3)]] =-1/(8u^8 v^9 w^6) ans..